The demand is given by P = 20 – 0.1Q, where P is the price and Q is the quantity demanded. The
monopolist’s total cost is C = 120 + 2Q + 0.05Q
. Find
a. profit-maximizing price and quantity. (2)
b. deadweight loss in the monopoly profit-maximizing equilibrium. (2)
1
Expert's answer
2020-04-13T08:58:26-0400

Solution:

a)


"MC=2+0.1Q"


"MC=p"

"2+0.1Q=20-0.1Q"

"Q=90"


"p=11"

b)


"20-0.1Q=0"


"Q=200"


"0.5 \times (200-90) \times 11=650"